# Hierarchical Procedural Framework for Low-latency Robot-Assisted Hand-Object Interaction

### v2

@@ 5.3.2 Adaptive Orientation
The robot handles environmental and human-induced disturbances in real-world implementations. Figure 5b shows the robot’s Z-axis translational and X-axis angular displacement trajectories. Over 30 seconds, the human bends and then lifts the index finger. The robot maintains its grip on the object (ring) and adjusts its TCP orientation in terms of vertical height and pointing angle. The robot achieves a maximum displacement of $\Delta Z$ of 74.23 mm and a rotational angle of 18.0 degrees. These adjustments enable the robot to maintain a suitable position for dressing, validating the effectiveness of the adaptive controller for action alignment discussed in Section 4.3. The error between the predicted TCP and the actual translational TCP is shown in Figure 5c, where the errors fluctuate around zero over time. The maximum absolute translational error is 2.48 mm, while the maximum absolute rotational error reaches 1.37 degrees.
@@ 5.4 Robot-assisted Ring Wearing
Figure 5d presents the results of the robot-assisted ring-wearing task (see supporting Video S3). The first and second rows of curves represent the TCP’s motions along the Y and Z axes, respectively, while the third row shows the angular displacement around the x-axis, corresponding to the ground coordinate system shown in Figure 3. In the open-loop hierarchy, the robot first performs a pre-programmed action to pick up the ring. It is then controlled by the MPM model, enabling the human operator to guide the end-effector’s approach using the ’Come’ hand pose. Once the operator specifies the desired location, they activate the ’Ring’ motion primitive, shifting the robot from passive control to active cooperation. From 17.6 seconds onward, the robot operates within the closed-loop hierarchy, continuously adjusting its movement towards the hand and dynamically compensating for any perturbations caused by the human hand. The dashed blue curve represents the motion of the CV-based hand pose estimation. As the human adjusts their finger position or location, the robot adapts in real-time, ensuring the ring remains properly oriented until it is successfully delivered. At 37.5 seconds, the robot completes the ring-wearing task by releasing the ring, thereby demonstrating the effectiveness of the robot in real-world applications such as robot-assisted dressing.
@@ 6 CONCLUSION
We proposed a hierarchical procedural framework for robot-assisted hand-object interaction. We presented an open-loop hierarchy that leverages the learning-based state estimation for real-time 3D construction of hand pose. An LSTM-based model has been trained to learn human hand motions. This is used as the motion primitives, which are translated into pre-defined robotic actions in a tele-manipulation manner. Regarding the low-level coordination hierarchy, we have designed a closed-loop controller that enables the spatial alignment between the end-effector and the fingering, ensuring that the object is manipulated collaboratively. Experimental tests have shown that the system exhibits a time delay of $\leq 0.3$ seconds, a translational error of $\leq 2.5$ mm, and a rotational error of $\leq 1.5$ degrees.

### v3

@@ 5.3.2 Adaptive Orientation
The robot handles environmental and human-induced disturbances in real-world implementations. Figure 5(b) shows the robot’s Z-axis translational and X-axis angular displacement trajectories. Over 30 seconds, the human bends and then lifts the index finger. The robot maintains its grip on the object (ring) and adjusts its TCP orientation in terms of vertical height and pointing angle. The robot achieves a maximum displacement of $\Delta Z$ of 74.23 mm and a rotational angle of 18.0 degrees. These adjustments enable the robot to maintain a suitable position for dressing, validating the effectiveness of the adaptive controller for action alignment discussed in Section 4.3. The error between the predicted TCP and the actual translational TCP is shown in Figure 5(c), where the errors fluctuate around zero over time. The maximum absolute translational error is 2.48 mm, while the maximum absolute rotational error reaches 1.37 degrees.
@@ 5.4 Robot-assisted Ring Wearing
Figure 5(d) presents the results of the robot-assisted ring-wearing task (see supporting Video S3). The first and second rows of curves represent the TCP’s motions along the Y and Z axes, respectively, while the third row shows the angular displacement around the x-axis, measured with respect to the ground coordinate system shown in Figure 3. In the open-loop hierarchy, the robot first performs a pre-programmed action to pick up the ring. It is then controlled by the MPM model, enabling the human operator to guide the end-effector’s approach using the ’Come’ hand pose. Once the operator specifies the desired location, they activate the ’Ring’ motion primitive, shifting the robot from passive control to active cooperation. From 17.6 seconds onward, the robot operates within the closed-loop hierarchy, continuously adjusting its movement towards the hand and dynamically compensating for any perturbations caused by the human hand. The dashed blue curve represents the motion of the CV-based hand pose estimation. As the human adjusts their finger position or location, the robot adapts in real-time, ensuring the ring remains properly oriented until it is successfully delivered. At 37.5 seconds, the robot completes the ring-wearing task by releasing the ring, thereby demonstrating the effectiveness of the robot in real-world applications such as robot-assisted dressing.
@@ 6 CONCLUSION
In this paper, we developed a hierarchical procedural framework for low-latency robot-assisted hand-object interaction. We presented an open-loop hierarchy that leverages learning-based state estimation for real-time 3D reconstruction of hand pose. Meanwhile, a LSTM-based model translates these poses into motion primitives, which are then executed as pre-defined robotic actions in a tele-manipulation manner. Regarding the low-level coordination hierarchy, we have designed a closed-loop controller that enables the spatial alignment between the end-effector and the fingering, ensuring that the object is manipulated collaboratively. Experimental tests have shown that the system exhibits a time delay of $\leq 0.3$ seconds, a translational error of $\leq 2.5$ mm, and a rotational error of $\leq 1.5$ degrees.